# BlockPruner: Fine-grained Pruning for Large Language Models

### v3

@@ Abstract
With the rapid growth in the size and complexity of large language models (LLMs), the costs associated with their training and inference have escalated significantly. Research indicates that certain layers in LLMs harbor substantial redundancy, and pruning these layers has minimal impact on the overall performance. While various layer pruning methods have been developed based on this insight, they generally overlook the finer-grained redundancies within the layers themselves. In this paper, we delve deeper into the architecture of LLMs and demonstrate that finer-grained pruning can be achieved by targeting redundancies in multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. We propose a novel, training-free structured pruning approach called BlockPruner. Unlike existing layer pruning methods, BlockPruner segments each Transformer layer into MHA and MLP blocks. It then assesses the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning. We applied BlockPruner to LLMs of various sizes and architectures and validated its performance across a wide range of downstream tasks. Experimental results show that BlockPruner achieves more granular and effective pruning compared to state-of-the-art baselines. Our code is available at: https://github.com/MrGGLS/BlockPruner.
@@ 1 Introduction
Building on these findings, we argue that finer-grained pruning can be effectively implemented in LLMs. Therefore, we introduce BlockPruner, a novel, training-free structured pruning approach. Unlike existing methods that focus on entire layers, BlockPruner segments each Transformer layer into multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. It then evaluates the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning.
@@ 2 Related Work
Concurrent and independent of our research, FINERCUT (Zhang et al., 2024a) also presents a fine-grained block pruning algorithm. However, their study does not delve into the rationale behind treating Transformer layers as two distinct sublayers for pruning purposes. In contrast, we began by conducting preliminary experiments that unveiled the fine-grained block redundancy within Transformer models. This discovery led us to propose the concept of minimal residual blocks. Furthermore, we explored how pruning various types of blocks affects the model’s performance. Furthermore, FINERCUT evaluates block importance by comparing the similarity between the output logits of the initial and pruned models. However, this metric may not adequately ensure that the pruned model generates coherent and semantically correct text, as it overlooks semantic information. In contrast, our approach measures block importance using the perplexity of the pruned model, which better reflects the fluency and quality of its output.
@@ 3 Methodology
The proposed fine-grained block pruning method (BlockPruner) is depicted in Figure 3. It begins by decomposing each Transformer layer into two minimal residual blocks (§3.1). We then evaluate the importance of each block using our proposed block importance metric (§3.2). Finally, we iteratively prune the block with the lowest importance (§3.3).
@@ 3.1 Minimal Residual Block
Equations (1) and (2) indicate that both types of residual blocks can be abstracted into a same computational formula. Therefore, intuitively treating them as independent layers for pruning appears reasonable, which will be further validated by our subsequent experimental results.
@@ 3.2 Block Importance
While previous layer pruning methods (Men et al., 2024; Chen et al., 2024) rely solely on the similarity between layer inputs and outputs to measure layer importance, we contend that this approach considers only the local influence of the layer while neglecting its role in the overall model’s output. To address the drawback, we introduce perplexity as a measure of block importance. Specifically, we determine the importance score of each block by masking it and then computing the perplexity of the new model on a given dataset. Intuitively, a block with the lowest importance score indicates that its removal results in minimal performance degradation. This method more effectively captures each block’s overall impact on the model’s performance, thereby more accurately reflecting its significance.
@@ 3.3 Iterative Search for Block Pruning
Unlike existing layer pruning techniques, which indiscriminately remove entire Transformer layers, we propose a novel fine-grained pruning strategy. This strategy selectively prunes multi-head attention (MHA) or multi-layer perceptron (MLP) blocks based on their defined importance. By employing this finer-grained pruning approach, we aim to better preserve the critical components and capabilities of the model while aggressively removing the less significant blocks.
For an LLM $\mathcal{M}$ with $L$ layers, we first divide them into $2L$ blocks, consisting of multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. Then, we perform iterative pruning search on a calibration dataset $\mathcal{C}$ to sequentially prune $K$ blocks. The steps are outlined as follows:
@@ 4.1 Experimental Setups
Table 1: Zero-shot downstream task performance of various models using different pruning methods. “Dense” represents the original, unpruned models. “PPL” means the perplexity on Wikitext2. All evaluations are conducted using the same configuration to ensure comparability.
We compare our method with several state-of-the-art structured pruning methods. The specific baseline methods include SliceGPT (Ashkboos et al., 2024), LaCo (Yang et al., 2024), ShortGPT (Men et al., 2024), and Relative Magnitude (Samragh et al., 2023; Men et al., 2024). SliceGPT achieves pruning by removing rows or columns corresponding to smaller principal components in the weight matrix. LaCo merges model layers from deep to shallow, using model output representations to calculate thresholds to avoid over-merging. ShortGPT eliminates redundant layers by calculating Block Influence. Relative Magnitude (RM) uses $||\frac{f(x)}{x+f(x)}||$ as an importance metric for layers, where $f(.)$ represents the non-residual part of the Transformer layer, and employs the same pruning method as ShortGPT. For SliceGPT, we used the official implementation [^2] As SliceGPT’s official code does not support Baichuan2 and Qwen1.5, we only employ it on the Llama2 series models.. For LaCo, we implemented it based on their code and controlled the number of pruned layers by adjusting the merging threshold. For ShortGPT and RM, we reproduced the results based on their paper descriptions.
@@
In our main experiment, we utilize the Alpaca dataset (Taori et al., 2023) to calculate importance scores. For our method, we employ only 256 samples to compute perplexity, and we discuss the influence of varying sample sizes in Section 5.4. Moreover, we observe that ShortGPT and Relative Magnitude are not sensitive to different numbers of samples. The effect of sample size on ShortGPT and Relative Magnitude is detailed in Appendix H. Nevertheless, we used the same number of samples as our method for consistency. All experiments in this study are conducted using two RTX 4090 GPUs.
@@ 4.2 Main Results
Previous structured pruning methods typically prune less than 30% of the parameters. Therefore, in our main experiments, we controlled the pruning ratio within this range. Since it is challenging to achieve identical pruning ratios across different methods and models, we select the closest available pruning ratios for comparison. The experimental results are presented in Table 1.
As shown in the results, our BlockPruner method significantly outperforms previous structured pruning baselines in terms of average performance and achieves the best results across most benchmarks, even though the pruning ratios in our method are slightly higher than that of baselines. We also observe that Llama2-13B maintains better performance at higher pruning ratios compared to Llama2-7B, with Baichuan2 and Qwen1.5 exhibiting similar behavior. This suggests that as the model scale grows, so does the number of redundant blocks, allowing for more pruning space.
Furthermore, it’s noteworthy that models exhibiting lower perplexity on the Wikitext2 dataset generally outperform those with higher perplexity on the same dataset. This underscores the potential of perplexity as a metric reflecting model performance. Notably, despite our method conducting pruning searches on the Alpaca dataset, it achieves lower perplexity on the Wikitext2 dataset.
Finally, we observe that while approaches such as ShortGPT and Relative Magnitude result in a significant decline in model performance across different tasks, BlockPruner stands out by avoiding such drastic reductions. This suggests that our proposed block pruning method effectively mitigates performance degradation during the pruning process. Due to space constraints, we have moved the details of pruning baselines and comparisons across various pruning ratios to Appendix I.
@@ 5.1 Ablation Study
To assess the influence of various key operations within the proposed pruning algorithm on its performance, we undertake a thorough ablation study across six models. In particular, we first drop the iterative search procedure and directly remove blocks with the lowest importance scores. Then, we substitute the fine-grained block pruning with a coarser-grained layer pruning approach. The results of these experiments are shown in Table 2.
@@
Furthermore, pruning at the layer level rather than the fine-grained block level yields less robust performance. This observation indicates that the model contains fine-grained redundancies, and segmenting layers into smaller blocks for pruning allows for more efficient removal of this redundancy, thereby better preserving the model’s capabilities.
@@ 5.2 Redundancies Between MHA and MLP
Before reaching a pruning ratio of 17%, pruning only the MHA blocks results in less performance loss compared to pruning MLP blocks and even matches the performance of mixed pruning. This indicates that MHA modules in LLMs may possess greater redundancy than initially anticipated, whereas MLP modules are relatively less redundant. However, when the pruning ratio surpasses 17%, further pruning of MHA blocks leads to a sharp decline in performance. This trend suggests that as pruning advances, the redundant MHA blocks are progressively removed, leaving only the crucial MHA blocks. Moreover, in the larger model, the sharp decline in performance occurs at higher pruning ratios. This observation is consistent with our previous findings, suggesting that larger models contain more redundant blocks. Such redundancy may stem from factors like insufficient training, resulting in higher initial redundancy.
Figure 4: The impact of pruning MHA and MLP individually on model performance. “MHA&MLP” represents the original BlockPruner algorithm.
@@
The results in Figure 5 (left) for both models reveal a consistent tendency to initially remove only MHA blocks. As the pruning process progresses and more blocks are removed, the proportion of MHA blocks being pruned follows a zigzag downward trend. Notably, the curve for Llama2-13B shifts to the right compared to Llama2-7B, suggesting that the larger model contains more redundant MHA blocks. This is further emphasized in Figure 5 (right), where, at the same pruning ratio, Llama2-13B prunes more MHA blocks than Llama2-7B. Additionally, given that our pruning method tends to remove more MHA blocks at equivalent pruning ratios, it can significantly reduce the usage of the key-value (KV) cache (Pope et al., 2023) in MHA, which potentially accelerate the inference process. To validate this, we also conducted a comparison of the inference speed among various models obtained through different pruning methods, with the results detailed in Appendix E.
@@ 5.3 Perplexity for Block Redundancy
However, as indicated in the ablation study, using perplexity without the iterative search procedure leads to a significant decline in performance. This suggests that perplexity alone is not an effective block importance metric. Instead, it is better suited for dynamic pruning algorithms that offer greater flexibility compared to static algorithms.
@@
In summary, perplexity leverages global information to effectively measure block redundancy, especially when used with a dynamic pruning strategy. This combination captures the complex interactions among blocks. In contrast, local metrics like Block Importance (BI) and Relative Magnitude (RM) are useful in specific scenarios but don’t reflect the overall contribution of blocks to the model, particularly at higher pruning ratios.
@@ 6 Conclusion
In this work, we introduce BlockPruner, a novel structured pruning approach for efficiently pruning large language models (LLMs). BlockPruner decomposes Transformer layers into two minimal residual blocks and employs a new block importance metric along with a pruning search algorithm to iteratively remove redundant blocks. Extensive experiments across various models show that our method outperforms other baselines in post-pruning performance. Our findings highlight the presence of fine-grained block redundancy in LLMs and reveal significant differences in redundancy levels among different block types. We hope our work contributes to a deeper understanding of the importance of various blocks within LLMs.
7 Limitations
@@ B.1 Pruning Datasets
Alpaca
is a general instruction-following dataset containing 52,000 questions. Each sample comprises three fields: instruction, input, and response. We selected 10% of the dataset and utilized 256 samples for the main experiments. Perplexity calculation was performed uniformly across all text in the samples without differentiation between fields.
@@ B.2 Evaluation Datasets
Wikitext-2
is a collection of over 100 million tokens extracted from verified Good and Featured articles on Wikipedia. This dataset is commonly used to measure the quality of a model’s text generation. We employed samples from the pre-split test set for calculating perplexity.
PIQA
is a dataset designed to evaluate natural language models’ understanding of physical commonsense. It employs a multiple-choice format where, given a goal and two possible solutions, the model must select the most appropriate solution.
WinoGrande
is an extensive dataset to evaluate models’ commonsense reasoning capabilities. It comprises 44,000 questions. The dataset features fill-in-the-blank tasks with binary options, aiming to select the correct option for a given sentence that requires commonsense reasoning.
HellaSwag
is also a dataset designed to assess models’ commonsense reasoning abilities, specifically to highlight the limitations of current models in handling commonsense natural language reasoning tasks. Despite being trivial for humans (with >95% accuracy), the dataset presents significant difficulties for models. The evaluation is conducted using four-way multiple-choice questions.
ARC
dataset comprises 7,787 multiple-choice science exam questions sourced from various origins. Each question typically offers four answer options. These questions are categorized into two distinct difficulty sets: 2,590 questions for Challenge Set and 5,197 for Easy Set.

### v4

@@ Abstract
With the rapid growth in the size and complexity of large language models (LLMs), the costs associated with their training and inference have escalated significantly. Research indicates that certain layers in LLMs harbor substantial redundancy, and pruning these layers has minimal impact on the overall performance. While various layer pruning methods have been developed based on this insight, they generally overlook the finer-grained redundancies within the layers themselves. In this paper, we delve deeper into the architecture of LLMs and demonstrate that finer-grained pruning can be achieved by targeting redundancies in multi-head attention (MHA) and multi-layer perceptron (MLP) blocks. We propose a novel, training-free structured pruning approach called BlockPruner. Unlike existing layer pruning methods, BlockPruner segments each Transformer layer into MHA and MLP blocks. It then assesses the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning. We applied BlockPruner to LLMs of various sizes and architectures and validated its performance across a wide range of downstream tasks. Experimental results show that BlockPruner achieves more granular and effective pruning compared to state-of-the-art baselines.
@@ 1 Introduction
Building on these findings, we argue that finer-grained pruning can be effectively implemented in LLMs. Therefore, we introduce BlockPruner, a novel, training-free structured pruning approach. Unlike existing methods that focus on entire layers, BlockPruner segments each Transformer layer into MHA and MLP blocks. It then evaluates the importance of these blocks using perplexity measures and applies a heuristic search for iterative pruning.
@@ 2 Related Work
Concurrent and independent of our research, FINERCUT (Zhang et al., 2024a) also presents a fine-grained block pruning algorithm. However, their study does not delve into the rationale behind treating Transformer layers as two distinct sublayers for pruning purposes. In contrast, we began by conducting preliminary experiments that unveiled the fine-grained block redundancy within Transformer models. This discovery led us to propose the concept of minimal residual blocks. Additionally, we explored how pruning different types of blocks impacts model performance. While FINERCUT assesses block importance by comparing the similarity between the output logits of the original and pruned models, this metric may fall short in ensuring that the pruned model produces coherent and semantically meaningful text, as it disregards semantic nuances. In our approach, we evaluate block importance using the perplexity of the pruned model, a metric that more effectively captures the fluency and quality of its outputs. To further support our perspective, we present a detailed comparison of these two metrics in Appendix D.
@@ 3 Methodology
The proposed fine-grained block pruning method (BlockPruner) is depicted in Figure 3. It begins by decomposing each Transformer layer into two minimal residual blocks (§3.1). We then evaluate the importance of each block by leveraging perplexity for our iterative block pruning framework (§3.2). Finally, we iteratively prune the block with the lowest importance (§3.3).
@@ 3.1 Minimal Residual Block
Equations (1) and (2) indicate that both types of residual blocks can be abstracted into the same computational formula. Hence, we argue that treating MLP and MHA blocks as the minimal units for pruning is a reasonable choice, which is substantiated by our subsequent experimental results.
@@ 3.2 Block Importance
While previous layer pruning methods (Men et al., 2024; Chen et al., 2024) rely solely on the similarity between layer inputs and outputs to measure layer importance, we argue that this approach overlooks the layer’s contribution to the overall model performance, while our metric considers its broader impact on the final output. To address the drawback, we introduce perplexity as a measure of block importance. Specifically, we determine the importance score of each block by masking it and then computing the perplexity of the new model on a given dataset. Intuitively, a block with the lowest importance score indicates that its removal results in minimal performance degradation. This method more effectively captures each block’s overall impact on the model’s performance, thereby more accurately reflecting its significance.
@@ 3.3 Iterative Search for Block Pruning
Unlike existing layer pruning techniques, which indiscriminately remove entire Transformer layers, we propose a novel fine-grained pruning strategy. This strategy selectively prunes MHA or MLP blocks based on their defined importance. By employing this finer-grained pruning approach, we aim to better preserve the critical components and capabilities of the model while aggressively removing the less significant blocks.
For an LLM $\mathcal{M}$ with $L$ layers, we first divide them into $2L$ blocks, consisting of MHA and MLP blocks. Then, we perform iterative pruning search on a calibration dataset $\mathcal{C}$ to sequentially prune $K$ blocks. The steps are outlined as follows:
@@ 4.1 Experimental Setups
Table 1: Zero-shot downstream task performance of various models using different pruning methods. “Dense” represents the original, unpruned models. “PPL” means the perplexity on Wikitext2.
We compare our method with several state-of-the-art structured pruning methods. The specific baseline methods include SliceGPT (Ashkboos et al., 2024), LaCo (Yang et al., 2024), ShortGPT (Men et al., 2024), and Relative Magnitude (Samragh et al., 2023; Men et al., 2024). SliceGPT achieves pruning by removing rows or columns corresponding to smaller principal components in the weight matrix. LaCo merges model layers from deep to shallow, using model output representations to calculate thresholds to avoid over-merging. ShortGPT eliminates redundant layers by calculating Block Influence. Relative Magnitude (RM) uses $||\frac{f(x)}{x+f(x)}||$ as an importance metric for layers, where $f(.)$ represents the non-residual part of the Transformer layer, and employs the same pruning method as ShortGPT. For SliceGPT, we used the official implementation [^2] As SliceGPT’s official code does not support Baichuan2 and Qwen1.5, we only employ it on the Llama2 series models.. For LaCo, we implemented it based on their code and controlled the number of pruned layers by adjusting the merging threshold. For ShortGPT and RM, we reproduced the results based on their manuscripts. More detailed implementation information is provided in Appendix A.
@@
In our main experiment, we utilize the Alpaca dataset (Taori et al., 2023) to calculate importance scores. For our method, we employ only 256 samples to compute perplexity, and we discuss the influence of varying sample sizes in Section 5.4. To ensure consistency, we use the same number of samples for ShortGPT and Relative Magnitude methods as shown in Appendix A. Moreover, the effect of sample size on ShortGPT and Relative Magnitude is detailed in Appendix I. All experiments are conducted on two RTX 4090 GPUs, and the execution times for different methods are reported in Appendix G.
@@ 4.2 Main Results
Prior studies (Yang et al., 2024; Ashkboos et al., 2024) have generally constrained the pruning ratio to approximately 25%. In line with these studies, we also restricted the pruning ratio to this range in our main experiments. Since it is challenging to achieve identical pruning ratios across different methods and models, we select the closest available pruning ratios for comparison.
As shown in Table 1, our BlockPruner method significantly outperforms previous structured pruning baselines in terms of average performance and achieves the best results across most benchmarks, even though the pruning ratios in our method are slightly higher than that of baselines. We also observe that Llama2-13B maintains better performance at higher pruning ratios compared to Llama2-7B, with Baichuan2 and Qwen1.5 exhibiting similar behavior. This suggests that as the model scale grows, so does the number of redundant blocks, allowing for more pruning space.
Furthermore, it’s noteworthy that models with lower perplexity on the Wikitext2 dataset tend to perform better, which highlights the correlation between perplexity and model effectiveness. This further supports the validity of perplexity as a reliable metric for evaluating model performance. Remarkably, although our method performs pruning searches on the Alpaca dataset, it achieves lower perplexity on the Wikitext2 dataset.
Finally, we observe that while approaches such as ShortGPT and Relative Magnitude result in a significant decline in model performance across different tasks, BlockPruner stands out by avoiding such drastic reductions. This suggests that our proposed block pruning method effectively mitigates performance degradation during the pruning process. Due to space constraints, we have moved the details of pruning baselines and comparisons across various pruning ratios to Appendix L.
@@ 5.1 Ablation Study
To assess the influence of various key operations within the proposed pruning algorithm on its performance, we undertake a thorough ablation study across six models. In particular, we first remove all blocks with the lowest importance scores at once, without the iterative search procedure. Then, we substitute the fine-grained block pruning with a coarser-grained layer pruning approach. The results of these experiments are shown in Table 2.
@@
Furthermore, pruning at the layer level rather than the block level yields less robust performance. This observation indicates that the model contains fine-grained redundancies, and segmenting layers into smaller blocks for pruning allows for more efficient removal of this redundancy, thereby better preserving the model’s capabilities. Additionally, we provide ablation experiments at higher sparsity levels, with results presented in Appendix E.
@@ 5.2 Redundancies Between MHA and MLP
Before reaching a pruning ratio of 17%, pruning only the MHA blocks results in less performance loss compared to pruning MLP blocks and even matches the performance of mixed pruning. This indicates that MHA modules in LLMs may possess greater redundancy than initially anticipated, whereas MLP modules are relatively less redundant. However, when the pruning ratio surpasses 17%, further pruning of MHA blocks leads to a sharp decline in performance. This trend suggests that as pruning advances, the redundant MHA blocks are progressively removed, leaving only the crucial MHA blocks. Moreover, in the larger model, the sharp decline in performance occurs at higher pruning ratios, which is consistent with the finding that larger models contain more redundant blocks. Such redundancy may stem from factors like insufficient training, resulting in higher initial redundancy.
Figure 4: The impact of pruning MHA and MLP individually with different pruning ratios on model performance. “MHA&MLP” represents the original BlockPruner algorithm. Results show that MHA modules in LLMs are more redundant than MLP modules.
@@
The results in Figure 5 (left) for both models reveal a consistent tendency to initially remove only MHA blocks. As the pruning process progresses and more blocks are removed, the proportion of MHA blocks being pruned follows a zigzag downward trend. Notably, the curve for Llama2-13B shifts to the right compared to Llama2-7B, suggesting that the larger model contains more redundant MHA blocks. This is further emphasized in Figure 5 (right), where, at the same pruning ratio, Llama2-13B prunes more MHA blocks than Llama2-7B. Additionally, given that our pruning method tends to remove more MHA blocks at equivalent pruning ratios, it can significantly reduce the usage of the key-value (KV) cache (Pope et al., 2023) in MHA, which potentially accelerate the inference process. To validate this, we also conducted a comparison of the inference speed among various models obtained through different pruning methods, with the results detailed in Appendix F.
@@ 5.3 Perplexity for Block Redundancy
However, as indicated in the ablation study, using perplexity without the iterative search procedure leads to a significant decline in performance. This suggests that while perplexity alone may not be a strong block importance metric, our iterative search method allows for a more effective use of it.
@@
In summary, perplexity leverages global information to effectively measure block redundancy, especially when used with a dynamic pruning strategy. This combination captures the complex interactions among blocks. In contrast, local metrics like BI and RM are useful in specific scenarios but don’t reflect the overall contribution of blocks to the model, particularly at higher pruning ratios.
@@ 6 Conclusion
In this work, we introduce BlockPruner, a novel structured pruning approach for efficiently pruning LLMs. BlockPruner decomposes Transformer layers into two minimal residual blocks and leverages a block importance metric based on perplexity in conjunction with an iterative pruning search algorithm, where the two components work together to progressively eliminate redundant blocks. Extensive experiments across various models show that our method outperforms other baselines in post-pruning performance. Our findings uncover fine-grained block redundancy in LLMs, highlighting significant differences in redundancy levels across different block types. We hope our work contributes to a deeper understanding of the importance of different blocks within LLMs.
Limitations
@@
Ethics Statement
The aim of this study is to provide a generalizable pruning method for large language models. All models and datasets used in our experiments are publicly accessible and do not contain any private information. We strictly adhere to the usage policies of these resources and utilize them solely for research purposes.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62176270) and the Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515012832).
@@ B.1 Pruning Datasets
Alpaca (Taori et al., 2023) is a general instruction-following dataset containing 52,000 questions. Each sample comprises three fields: instruction, input, and response. We selected 10% of the dataset and utilized 256 samples for the main experiments. Perplexity calculation was performed uniformly across all text in the samples without differentiation between fields.
@@ B.2 Evaluation Datasets
Wikitext-2 (Merity et al., 2016) is a collection of over 100 million tokens extracted from verified Good and Featured articles on Wikipedia. This dataset is commonly used to measure the quality of a model’s text generation. We employed samples from the pre-split test set for calculating perplexity.
PIQA (Bisk et al., 2020) is a dataset designed to evaluate natural language models’ understanding of physical commonsense. It employs a multiple-choice format where the model selects the most appropriate solution from two options given a goal.
WinoGrande (Sakaguchi et al., 2021) is an extensive dataset to evaluate models’ commonsense reasoning capabilities. It comprises 44,000 questions. The dataset features fill-in-the-blank tasks with binary options, aiming to select the correct option for a given sentence that requires commonsense reasoning.
HellaSwag (Zellers et al., 2019) is also a dataset designed to assess models’ commonsense reasoning abilities, specifically to highlight the limitations of current models in handling commonsense natural language reasoning tasks. Despite being trivial for humans (with >95% accuracy), the dataset presents significant difficulties for models. The evaluation is conducted using four-way multiple-choice questions.
ARC (Clark et al., 2018) dataset comprises 7,787 multiple-choice science exam questions sourced from various origins. Each question typically offers four answer options. These questions are categorized into two distinct difficulty sets: 2,590 questions for Challenge Set and 5,197 for Easy Set.